a 300 kg has passenger with an unknown mass. From rest, the elevator accelerates upward at 1.2m/s/s. The cable supporting the elevator exerts and upward force of 8000 N. What is the maximum mass of the passenger that the elevator can carry?

1
Expert's answer
2022-03-16T09:42:08-0400

The tension in the cable

"T=m(g+a)"

Hence, the net mass of system (elevator and passengers) is equal

"m=T/(g+a)\\
=8000/(9.8+1.2)=727\:\rm kg"

The limit mass of passengers

"m_p=727-300=427\:\rm kg"